# CoTAL: Human-in-the-Loop Prompt Engineering, Chain-of-Thought Reasoning, and Active Learning for Generalizable Formative Assessment Scoring

### v2

@@ V Analyzing RQ1: CoTAL Generalizability Across Multiple Connected Domains
V-A Rules Task
Performance results for the Rules Task comparing CoTAL and the Baseline are presented in Table IV.
@@ V-A Rules Task
TABLE IV: Performance results for CoTAL applied to the Rules Task relative to the Baseline in terms of Cohen’s QWK. Each “R” corresponds to a different subscore for the Rules Task and is explained in Table I. Total Score compares the LLM’s prediction of the total score (i.e., the sum of all 9 subscores) to the human label. For each metric, the best-performing implementation is in boldface.
The Rules Task Baseline implementation resulted in an average QWK of 0.826 across all subscores. Applying CoTAL resulted in an average QWK of 0.916, which represents an average increase of 0.090 (10.9% ) over the Baseline while using CoTAL. By itself, the Baseline implementation provided good results with 7/9 Baseline subscores showing “strong” agreement [40]. Incorrect predictions were evenly split between false positives and negatives (7 and 6, respectively). The Baseline underscored the Total Score by exactly one point in 7 of the 32 test set instances because it failed to properly add the predicted subscores (e.g., the LLM awarded 8 subscore points but a Total Score of 7). Interestingly, this “one under” issue was not observed in previous work [12]. We hypothesize that a large number of subscores (nine for the Rules Task) may cause issues during inference due to the exponential nature of combinatorics, resulting in $5122^{9}=$ 512 scoring combinations. Further research is necessary to evaluate this hypothesis empirically. The LLM Baseline incorrectly predicted 17/32 Total Scores overall (12 of which were underscores).
CoTAL improved results, with every individual subscore and Total Score (except R5) beating or tying the Baseline. CoTAL reduced false positives in individual subscores from 7 to 1 while only increasing false negatives by 1. This was explicitly addressed during Active Learning, which favored false positives to false negatives by a 2:1 ratio. R5 was the only subscore with more than one incorrect prediction with CoTAL, while the Baseline had 4/9 subscores with multiple incorrect predictions. CoTAL resulted in 10 incorrectly predicted Total Scores, compared to 17 for the Baseline implementation, with 8 instances of underscoring and 2 of overscoring. Only 2 of the 10 incorrectly predicted Total Scores were off by two or more points, compared to 5 in the Baseline approach. CoTAL also reduced the Baseline’s “one under” Total Score addition errors from 7 with the Baseline to 2 using CoTAL.
In addition to accurate scoring, the LLM provided accurate feedback for the Rules Task, citing relevant evidence, tying it to specific rubric components, and justifying its scoring decisions like a human scorer. The following example highlights a successful LLM generation for R4 in the Rules Task.
@@
As the example shows, the student correctly identified R4 and set absorption to the absorption limit. Interestingly, the LLM cited the correct portion of the student’s response and acknowledged the rubric’s directive but did not award a point. This is a clear case of logical inconsistency (i.e., LLM outputs that exhibit logical contradictions or are inconsistent with their CoT reasoning chains [42]), as the LLM disregards acknowledged information when scoring. This demonstrates the LLM can produce inaccurate scores even when the correct score is obvious to a human. While rare, this also highlights that the LLM can arrive at the wrong answer even when citing the correct portion of the student’s response.
Utilizing CoTAL for the Rules Task demonstrated the LLM’s proficiency in generating accurate scoring predictions and delivering valuable feedback when prompted effectively. This indicates that the LLM can serve as a robust tool for scoring and feedback, especially when CoTAL is employed in conjunction with clearly defined formative assessment questions and rubrics.
V-B Debugging Task
The performance comparison between CoTAL and the Baseline for the Debugging Task appears in Table V. For this task, the LLM needed access to the fictional student’s erroneous computational model (see Section III), which we distilled into textual form for inclusion in our prompt. Token limitations in GPT-4’s context window inhibited Active Learning in the Debugging Task, so the results only include the Response Scoring and Prompt Development components of CoTAL Phase II.
@@ V-B Debugging Task
TABLE V: CoTAL versus the Baseline Performance for the Debugging Task using Cohen’s QWK. Each “D” corresponds to a different subscore for the Debugging Task and is explained in Table II. Total Score compares the LLM’s prediction of the total score (i.e., the sum of all 5 subscores) to the human-assigned score. For each metric, the best-performing implementation is in boldface.
The Debugging Task Baseline implementation resulted in an average QWK of 0.567, while applying CoTAL generated an average QWK of 0.706, representing a 24.5% increase. Baseline performance for the Debugging Task was worse than the Rules Task, with only 2 of 5 subscores having QWKs $\geq 0.8$. This was due to the LLM’s tendency to generate false positives, which outnumbered false negatives 19 to 2. The LLM’s Total Score predictions did not suffer from the “one under” issue but did not align as well with the human scorers (QWK$=0.561$). The LLM correctly predicted the Total Score in 23/33 instances, with all 10 incorrect predictions being overscoring due to false positives. Six of these incorrect predictions were overscored by 2 or more points.
CoTAL produced significantly better results, even without Active Learning. All Debugging Task subscores and Total Scores improved. The false positive to false negative ratio dropped from 19:2 with the Baseline to 2:1 using CoTAL. The severity of the imbalance was reduced, and the number of incorrect predictions decreased, except for D5. The LLM produced five fewer incorrect predictions across subscores with CoTAL than with the Baseline. For Total Score, QWK rose by 0.218 (a 38.9% gain), and the number of instances incorrectly scored by two or more points dropped from 6 to 2, indicating better alignment with human scorers. The LLM provided feedback linked to the student’s response, rubric, prompt guidelines, and computational model using CoTAL. An example of successful LLM feedback for D1 is shown below.
@@
In the example, the LLM isolates the specific portion of the student’s response that corresponds to D1. The LLM cites the student, explains why the response earns a point according to the rubric and provides the correct score (1). This example highlights the LLM’s ability to identify relevant portions of the student’s response, connect it to the rubric, and assign the correct score. The LLM often referenced other prompt components, such as guidelines and the computational model, citing specific line numbers and code blocks to clarify its scoring decisions. This demonstrates that CoTAL enables the LLM to discern meaning from textual representations of block-based code, even without Active Learning.
Although the LLM effectively interpreted the computational model for scoring and generating feedback, it sometimes struggled with the extra context. As a result, the LLM occasionally had trouble tracking all relevant information. The example below depicts an instance where the LLM incorrectly awarded a point for D3.
@@
In this example, the LLM correctly identified the “third rule” (from the Rules Task) as the “greater than” condition but incorrectly used it to award a point for D3, which belongs to the “less than” condition. The LLM linked the “greater than” condition to the “set [Absorption (inch)] to [Absorption Limit (inch)]” code block on Line 8, though the “greater than” condition did not appear until Line 10. It then referenced the rubric for D3 and the “less than” condition before erroneously awarding a point for D3. This is a case of context inconsistency (i.e., LLM outputs that are unfaithful to the user’s provided context where the LLM contradicts a fact in the prompt [42]), where the LLM mixed up “rules,” corresponding “if statements,” and line numbers, demonstrating difficulties in tracking multiple pieces of information. This example also highlights the interconnectedness of our formative assessments, rubrics, and prompts, requiring domain knowledge from the Rules Task to assess and provide feedback for Debugging Task responses.
@@
V-C Engineering Task
Unlike the previous Rules and Debugging Tasks, the Engineering Task rubric did not have multiple subscores. Instead, students received a single score in the range 0 to 4 (inclusive) based on the rubric in Table III. QWK for the Baseline was 0.970, while QWK using CoTAL was 0.965. The Engineering Task Baseline resulted in four incorrect scoring predictions out of 32, and the QWK score of 0.97 was comparable to the agreement between the two humans during IRR (QWK = 0.975). This supports our previous hypothesis that the human scorer agreement during IRR may indicate the LLM’s initial alignment with human consensus [12].
CoTAL reduced incorrect predictions from 4 to 2, with QWK performance falling slightly by 0.005 due to one instance deviating from the human score by more than one point. This reveals two issues: (1) a decrease in QWK doesn’t imply reduced accuracy, as CoTAL achieved higher accuracy with fewer errors; and (2) prompt engineering may yield diminishing returns for tasks where the LLM performs well out-of-the-box. The performance improvement with CoTAL is attributed to Active Learning, where the LLM initially overlooked students’ discussions on engineering constraints. Students mentioned specific considerations like “cost” and “accessible squares,” but the LLM did not connect these to the rubric. After integrating a validation set instance into the prompt and using corrective CoT reasoning chains, the LLM’s accuracy improved.
Similar to the Rules and Debugging Tasks, the LLM assigned accurate scores. It provided helpful feedback using CoTAL to cite evidence from students’ responses and connected it with the rubric. For example, when a student answered “yes” in the Engineering Task (incorrect answer; 0 points), the LLM still offered a constructive critique:
Student: Yes. Morgan’s second design is better because its cheaper and has less runoff.
@@ V-C Engineering Task
This example illustrates that the LLM correctly awarded a score of 0, as the student responded “yes” to the assessment question. However, the LLM’s feedback also highlighted the student’s understanding of key design constraints—cost and runoff—even though no points were earned. This recognition is important: it surfaces evidence of conceptual understanding that might otherwise go unnoticed, helping teachers acknowledge student understanding and providing them with affirming, formative feedback despite an incorrect response. Without this level of explanation, educators might overlook the student’s grasp of domain concepts.
@@
We also noted instances of undesirable LLM behavior. Similar to the Rules and Debugging Tasks (as well as the Science Concepts and Reasoning Task from earlier research), the LLM can be deceived by misleading responses. In one case, a student stated, “Morgan needs to check how other amounts of rainfall affect her design,” for which the LLM awarded 3 points. However, to earn 3 points, students needed to identify the difference in rainfall between tests, not the amounts. This issue was noted during Response Scoring and extensively discussed by the research team. Despite emphasizing this in the rubric and prompt guidelines, the LLM still scored this answer incorrectly.
@@
Another issue was that the LLM sometimes included flawed reasoning in its scoring explanations. In one instance, a student responded: “No. Because she one has better cost and worse absorption, and the other has better absorption and worse cost.” The LLM cited the correct portion of the student’s response about engineering constraints, noting it “shows an understanding of the trade-offs between the Engineering Constraints”. Human scorers were clear that this response scored two points based on the rubric. However, the LLM incorrectly stated, “the student does not mention the different rainfall values or the specific Engineering Constraints by name, so the student cannot be awarded 2, 3, or 4 points.” This is an instruction inconsistency hallucination, where the LLM deviates from an explicit instruction provided by the user [42] (the prompt does not require students to identify constraints “by name” for credit). During Active Learning, we used CoT reasoning to show that responses mentioning trade-offs between absorption, runoff, and cost should receive 2 points. Despite recognizing the student’s understanding of the constraints, the LLM failed to award the correct score of 2 points.
@@
Overall, the LLM effectively scored student responses and provided clear rationales using CoTAL. Like the Rules Task, the Engineering Task showed the LLM’s ability to score and give feedback explicitly linked to the rubrics and student comprehension. This study and previous research typically used binary multi-label scoring, but the Engineering Task demonstrated the LLM’s effectiveness with a multi-class (5-way) scoring scheme. Although CoTAL resulted in a slight drop in QWK score and some hallucinations, it allowed the LLM to maintain near-perfect alignment with human scorers and explain scores accurately based on the rubric.
V-D Answering RQ1
@@
Overall, CoTAL generalized well across all tasks and domains, improving average QWK performance by 10.9% and 24.5% for the Rules and Debugging Tasks, respectively. In the Engineering Task, CoTAL predicted two fewer incorrect instances despite a slight QWK drop. Using CoTAL, the LLM agreed with researcher evaluations for 94.7% of the 550 answers across all three formative assessments’ test sets, resulting in errors on 29 of them, where the LLM either fabricated information or otherwise produced outputs that diverged from the researchers’ preferences. Hallucinations were context-dependent, stemming from individual misunderstandings rather than being domain-specific.
@@
Qualitatively, CoTAL produced LLM outputs that accurately explained scores by citing correct evidence from the student’s response and tying it to the rubrics. Although additional mechanisms are needed to reduce LLM hallucinations, CoTAL improved scoring accuracy and enabled LLMs to provide interpretable scores and explanations across multiple domains and assessments. Every formative assessment required its own context-specific prompt (e.g., including domain concepts, assessment questions, and rubrics), yet the same prompt engineering procedure was effective across science, computation, and engineering without methodological adjustments.
VI Analyzing RQ2: Educator and Student Feedback

### v3

@@ 5 Analyzing RQ1: CoTAL Generalizability Across Multiple Connected Domains
5.1 Rules Task
Performance results for the Rules Task comparing CoTAL and the Baseline are presented in Table 4. The Rules Task Baseline implementation resulted in an average QWK of 0.826 across each subscore. [^7] Among the metrics we report, the two aggregate measures include: (1) average subscore QWK, calculated as the mean of the QWK values across all rubric subscores (e.g., R1 to R9 for F2), i.e., $\frac{1}{n}\sum_{i=1}^{n}\text{QWK}(S_{i},\hat{S}_{i})$; and (2) Total Score QWK, computed as $\text{QWK}(\hat{T},\sum_{i=1}^{n}S_{i})$, where $\hat{T}$ is the LLM’s predicted Total Score, $\sum_{i=1}^{n}S_{i}$ is the sum of ground truth subscores, and $n$ is the number of subscores defined by the rubric. Applying CoTAL resulted in an average QWK of 0.916, which represents an average increase of 0.090 (10.9% ) over the Baseline while using CoTAL. By itself, the Baseline implementation provided good results with 7/9 Baseline subscores showing “strong” agreement [56]. Incorrect predictions were evenly split between false positives and negatives (7 and 6, respectively). The Baseline underscored the Total Score by exactly one point in 7 of the 32 test set instances because it failed to properly add the predicted subscores (e.g., the LLM awarded 8 subscore points but its predicted Total Score was 7). Interestingly, this “one under” issue was not observed in previous work [19]. We hypothesize that a large number of subscores (nine for the Rules Task) may cause issues during inference due to the exponential nature of combinatorics, resulting in $2^{9}=512$ scoring combinations. Further research is necessary to evaluate this hypothesis empirically (i.e., investigating why the LLM incorrectly summed its own subscore predictions so often). The LLM Baseline incorrectly predicted 17/32 Total Scores overall (12 of which were underscores).
CoTAL improved results, with every individual subscore and Total Score (except R5) beating or tying the Baseline. CoTAL reduced false positives in individual subscores from 7 to 1 while only increasing false negatives by 1. This was explicitly addressed during Active Learning, which favored false positives over false negatives by a 2:1 ratio. R5 was the only subscore that had more than one incorrect prediction with CoTAL, while the Baseline had 4/9 subscores with multiple incorrect predictions. CoTAL resulted in 10 incorrectly predicted Total Scores, compared to 17 for the Baseline implementation, with 8 instances of underscoring and 2 of overscoring. Only 2 of the 10 incorrectly predicted Total Scores were off by two or more points, compared to 5 in the Baseline approach. CoTAL also reduced the Baseline’s “one under” Total Score versus Baseline addition errors from 7 to 2.
In addition to accurate scoring, the LLM provided accurate explanations for the scores assigned to the Rules Task assessment responses, citing relevant evidence, tying it to specific rubric components, and justifying its scoring decisions like a human scorer. The following example highlights a successful LLM generation for R4 in the Rules Task.
@@ 5.1 Rules Task
Table 4: Performance results for CoTAL applied to the Rules Task relative to the Baseline in terms of Cohen’s QWK. Each “R” corresponds to a component subscore for the Rules Task and is explained in Table 1. Total Score compares the LLM’s prediction of the total score (i.e., the sum of all 9 subscores) to the human label. For each metric, the best-performing implementation is in boldface.
@@
As the example shows, the student correctly identified R4 and set absorption to the absorption limit. The LLM correctly cited the relevant portion of the student’s response, acknowledged the applicable rubric directive, yet still failed to award the appropriate point. This is a clear case of logical inconsistency, where the LLM output exhibits logical contradictions or is inconsistent with its CoT reasoning chains or cited evidence [58].
Using CoTAL for the Rules Task showcased the LLM’s ability to generate accurate scoring predictions and provide relevant feedback when prompted appropriately. This evaluation illustrates CoTAL’s intended purpose: to facilitate scoring and feedback generation based on well-defined formative assessment questions and rubrics. As a result, it allows for a systematic assessment of the LLM’s proficiency when given effective prompts.
5.2 Debugging Task
The performance comparison between CoTAL and the Baseline for the Debugging Task appears in Table 5. For this task, the LLM needed access to the fictional student’s erroneous computational model (see Section 3), which we distilled into textual form for inclusion in our prompt. Token limitations in GPT-4’s context window prohibited Active Learning in the Debugging Task, so the results only include the Response Scoring and Prompt Development components of CoTAL Phase II.
@@ 5.2 Debugging Task
Table 5: CoTAL versus the Baseline Performance for the Debugging Task using Cohen’s QWK. Each “D” corresponds to a subscore for the Debugging Task, which is explained in Table 2. Total Score compares the LLM’s prediction of the total score (i.e., the sum of all 5 subscores) to the human-assigned score. For each metric, the best-performing implementation is in boldface.
For the Debugging Task, the Baseline implementation achieved an average QWK of $0.567$, whereas CoTAL yielded an average QWK of $0.706$, a 24.5% improvement. Baseline performance lagged behind the Rules Task, with only 2 of 5 subscores attaining QWKs $\geq 0.8$, primarily due to a strong tendency toward false positives (FPs) that outnumbered false negatives (FNs) 19 to 2. Total Score predictions did not exhibit the “one under” pattern but aligned less well with human scoring (QWK=$0.561$). The LLM was correct in 23/33 cases, and all 10 errors were due to overscoring (i.e., FPs)—including six instances that were overscored by 2 or more points. Applying CoTAL reduced this imbalance and decreased incorrect predictions across subscores, except for D5. Compared to the Baseline, the LLM made five fewer subscore errors. For Total Score, QWK increased by 0.218 (a 38.9% gain), and the number of instances with incorrect scores of two or more points declined from 6 to 2, indicating improved alignment with human scorers.
Under CoTAL, the LLM’s feedback explicitly linked cited student text to rubric criteria, prompt guidelines, and the computational model. An example of successful LLM feedback for D1 is shown below.
@@
In the example, the LLM isolates the specific portion of the student’s response that corresponds to D1. The LLM cites the student, explains why the response earns a point according to the rubric, and provides the correct score (1). This example highlights the LLM’s ability to identify relevant portions of the student’s response, connect it to the rubric, and assign the correct score. The LLM often referenced other prompt components, such as guidelines and the computational model, citing specific line numbers and code blocks to clarify its scoring decisions. This demonstrates that CoTAL enables the LLM to discern meaning from textual representations of block-based code, even without Active Learning.
While the LLM generally interpreted the computational model correctly for scoring and feedback, the additional contextual information occasionally impeded information tracking. Consequently, it sometimes misapplied criteria, as illustrated by a case in which it incorrectly awarded a point for D3.
@@
In this case, the LLM correctly identified the “third rule” (from the Rules Task) as the “greater than” condition but erroneously used it to award a point for D3, which pertains to the “less than” condition. The LLM linked the “greater than” condition to the “set [Absorption (inch)] to [Absorption Limit (inch)]” code block on Line 8. However, the “greater than” condition did not appear until Line 10. It then cited the D3 rubric and the “less than” condition before incorrectly awarding credit. This constitutes context inconsistency—outputs unfaithful to the provided context in which the LLM contradicts a fact in the prompt [58]—arising from a mix-up of “rules,” corresponding “if statements,” and line numbers, and indicating difficulty tracking multiple information sources. The example also underscores the interconnectedness of our formative assessments, rubrics, and prompts, which require domain knowledge from the Rules Task to assess and provide feedback for Debugging Task responses.
@@
5.3 Engineering Task
Unlike the previous Rules and Debugging Tasks, the Engineering Task rubric did not have multiple subscores. Instead, students received a single score in the range 0 to 4 (inclusive) based on the rubric in Table 3. QWK for the Baseline was $0.97$, while QWK using CoTAL was $0.965$. The Engineering Task Baseline resulted in four incorrect scoring predictions out of 32, and the QWK score of $0.97$ was comparable to the agreement between the two humans during IRR (QWK =0.975=0.975). This supports our previous hypothesis that the human scorer agreement during IRR may indicate the LLM’s initial alignment with human consensus [19].
CoTAL reduced the number of incorrect predictions from 4 to 2. Although the QWK decreased marginally by $0.005$, this reduction was attributable to a single instance that deviated from the human-assigned score by more than one point. These results highlight two key observations: (1) a decrease in QWK does not necessarily indicate reduced accuracy, as CoTAL achieved higher accuracy with fewer errors; and (2) prompt engineering may yield diminishing returns for tasks where the LLM already performs well out of the box. CoTAL’s observed accuracy gain is attributable to Active Learning (Phase II, Step 3 of Figure 3). Students mentioned specific considerations like “cost” and “accessible squares,” but the LLM did not connect these to the rubric. After incorporating a validation-set instance into the prompt and introducing corrective CoT reasoning chains, the LLM’s accuracy improved.
Consistent with the Rules and Debugging Tasks, the LLM assigned accurate scores and, using CoTAL, provided constructive feedback by citing evidence from students’ responses and aligning it with the rubric. For example, when a student answered “yes” in the Engineering Task (incorrect answer; 0 points), the LLM nevertheless offered a constructive critique:
Student: Yes. Morgan’s second design is better because it’s cheaper and has less runoff.
@@ 5.3 Engineering Task
This example illustrates that the LLM correctly awarded a score of 0, as the student responded “yes” to the assessment question. However, the LLM’s feedback also highlighted the student’s understanding of key design constraints—cost and runoff—even though no points were earned. This recognition is important: it surfaces evidence of conceptual understanding that might otherwise go unnoticed if only a student’s score is considered. This helps teachers acknowledge student understanding and providing them with affirming, formative feedback despite an incorrect response. Without this level of explanation, teachers might overlook the student’s grasp of domain concepts.
@@
We also observed instances of undesirable LLM behavior. Consistent with the Rules and Debugging Tasks (and the Science Concepts and Reasoning Task reported in prior work), the LLM was susceptible to misleading responses. In one case, a student wrote, “Morgan needs to check how other amounts of rainfall affect her design,” which the LLM scored as 3 points. However, earning 3 points required identifying the difference in rainfall between tests, not the amounts. This issue emerged during Response Scoring and was extensively discussed by the research team. Despite emphasizing this distinction in both the rubric and the prompt guidelines, the LLM nevertheless assigned an incorrect score.
@@
A second issue concerned flawed reasoning in the LLM’s scoring explanations. For example, a student responded: “No. Because she one has better cost and worse absorption, and the other has better absorption and worse cost.” The LLM correctly cited the portion of the response referencing engineering constraints, noting it “shows an understanding of the trade-offs between the Engineering Constraints.” Human scorers consistently assigned this response two points according to the rubric. Nonetheless, the LLM incorrectly asserted, “the student does not mention the different rainfall values or the specific Engineering Constraints by name, so the student cannot be awarded 2, 3, or 4 points.” This constitutes an instruction inconsistency hallucination, wherein the LLM deviates from explicit user instructions [58]; notably, the prompt does not require students to identify constraints “by name” to receive credit. During Active Learning, we provided CoT reasoning demonstrating that responses referencing trade-offs among absorption, runoff, and cost warrant 2 points. Despite recognizing the student’s understanding of these constraints, the LLM failed to assign the correct score.
@@
Overall, the LLM effectively scored student responses and provided clear rationales using CoTAL. As in the Rules Task, the Engineering Task highlighted the LLM’s capacity to assign scores and deliver feedback explicitly aligned with the rubric and student understanding. Whereas this study and prior work primarily employed binary multi-label scoring, the Engineering Task further demonstrated the LLM’s effectiveness under a multi-class (5-way) scoring scheme. Although CoTAL produced a slight decline in QWK and occasional hallucinations, it enabled the LLM to maintain near-perfect alignment with human scorers and to justify scores accurately using the rubric.
5.4 Answering RQ1
@@
Overall, CoTAL generalized effectively across tasks and domains, increasing average subscore QWK by 10.9% and 24.5% for the Rules and Debugging Tasks, respectively. In the Engineering Task, CoTAL reduced the number of incorrect predictions by two despite a slight decrease in QWK. Using CoTAL, the LLM agreed with researcher evaluations on 94.7% of the 550 answers across the three formative assessments’ test sets, yielding errors on 29 instances in which the model either fabricated information or otherwise produced outputs that diverged from human preferences. These hallucinations were context-dependent, arising from individual misunderstandings rather than domain-specific weaknesses.
@@
Qualitatively, CoTAL generated outputs that accurately justified scores by citing appropriate evidence from student responses and explicitly linking this evidence to the rubrics. While additional mechanisms are needed to reduce hallucinations further, CoTAL improved scoring accuracy and enabled LLMs to provide interpretable scores and explanations across multiple domains and assessments. Each formative assessment required a context-specific prompt (e.g., domain concepts, assessment questions, and rubrics); nevertheless, the same prompt-engineering procedure was effective across science, computation, and engineering without methodological adjustments.
6 Analyzing RQ2: Teacher and Student Feedback